Parker Plastic, Inc., manufactures plastic mats to use with rolling office chairs. Its standard cost information for last year follows:

Standard Quantity Standard Price (Rate) Standard Unit Cost
Direct materials (plastic) 12 sq ft. $ 0.83 per sq. ft. $ 9.96
Direct labor 0.25 hr. $ 10.50 per hr. 2.62
Variable manufacturing overhead (based on direct labor hours) 0.25 hr. $ 2.20 per hr. 0.55
Fixed manufacturing overhead $345,800 ÷ 910,000 units) 0.38


Parker Plastic had the following actual results for the past year:

Number of units produced and sold 1,040,000
Number of square feet of plastic used 11,400,000
Cost of plastic purchased and used $ 9,120,000
Number of labor hours worked 308,000
Direct labor cost $ 3,449,600
Variable overhead cost $ 689,000
Fixed overhead cost $ 365,000

Calculate Parker Plastic’s direct materials price and quantity variances. (Do not round intermediate calculations. Indicate the effect of each variance by selecting favorable, unfavorable, or no effect (i.e., zero variance).)
Answer #1

Direct Material Price Variance = (Actual price - standard price) x actual quantity purchased
= ($0.80 - $0.83) x 11400000 = $342000 (F)
Actual Price = $9120000 / 11400000 = $0.80

Direct Material Quantity Variance = (Actual quantity - standard quantity) x Standard Price
= (11400000 - 12480000) x $0.83 = $896400 (F)
Standard Quantity = 1040000 x 12 = 12480000